OpenStudy (anonymous):

Can anyone tell me how to solve this equation f(x)=x^3-5X^2+7x+2 were the x values are 0,1,2,3,4,5 Thanks

OpenStudy (anonymous):

just plug in those values for x and you'll have the solution f(0) = 0^3 - 5(0)^2 + 7(0) + 2 = 2 and so on with the others

OpenStudy (anonymous):

Thanks for the help when I plug in the 1 I get a -15 is that the same thing you get?

OpenStudy (anonymous):

nope. for 1 i get: f(1)=(1)^3-5(1)^2+7(1)+2 f(1)= 1 - 5 + 7 + 2 f(1)= 5